OpenStudy (anonymous):

Can somebody explain how my solutions manual got y=o for solving the following equation? (3 - 5y)/ (2 + y) = (3 + 5y)/ (2 - y) I've tried inserting 0 for y...and it's not coming out right for me.

OpenStudy (anonymous):

if you replace y by 0 both sides are \[\frac{3}{2}\] yes. so that shows the answer is correct even if it doesn't explain how to get it

OpenStudy (anonymous):

i would try "cross multiplying' to get \[(2-y)(3-5x)=(2+y)(3+5x)\]

OpenStudy (anonymous):

I'm trying to get it to equal 3/2 on each side..feel free to go elsewhere on the sight, while I try and get this to work..

OpenStudy (anonymous):

light bulb came on, thanks..

OpenStudy (anonymous):

i have no idea why i put those x's there. let me write it correctly

OpenStudy (anonymous):

\[(2-y)(3-5y)=(2+y)(3+5y)\]

OpenStudy (anonymous):

multiply out and you will see that y = 0